# Semitoric Families on Pentagon Spaces

### v1

@@ Abstract.
Semitoric systems are a special type of 4-dimensional integrable system where one of the functions is the moment map of a Hamiltonian $S^{1}$-action. These systems were classified by Pelayo and Vũ Ngọc, but there are not many explicit examples. Recently, Le Floch and Palmer introduced semitoric transition families in which a singular point transitions between elliptic-elliptic type and focus-focus type as the parameter varies. In this paper, we construct semitoric transition families on pentagon spaces by interpolating two semitoric systems of toric type. More specifically, we show that there exists a unique singular point that changes from elliptic-elliptic type to focus-focus type and back to elliptic-elliptic type as the parameter varies.
@@ 1. Introduction
An integrable system on a $2n$-dimensional symplectic manifold $(M,\omega)$ is a map $F=(f_{1},\ldots,f_{n}):M\to{\mathbb{R}}^{n}$ such that $\{f_{i},f_{j}\}=0$ for all $i,j$ and $df_{1},\ldots,df_{n}$ are linearly independent almost everywhere. A semitoric system is a special kind of integrable system when $\dim M=4$. More specifically, an integrable system $(M,\omega,F=(J,H))$ is a semitoric system if $J$ is a proper map that generates an effective Hamiltonian $S^{1}$-action on $M$ and the momentum map $F$ only has non-degenerate singularities without hyperbolic type (see Definition 3.3). Semitoric systems were introduced by  [VuN07] following the work of [Sym03]. Pelayo and Vũ Ngọc [PVN09, PVN11] first classified semitoric systems with generic critical fibers. Following their work, Pelayo-Palmer-Tang [PPT24] generalized the classification to any semitoric systems. Semitoric systems have been intensively studied in the recent years. See [ADH20, ADH19, AHP25, LFP19] for studies of invariants of semitoric systems and [AH19, SVN17, DKL+23] for recent survey articles and additional references.
The coupled angular momenta system was the first example of a family of semitoric system known to us. It was first introduced in [SZ99] and studied in detail in [LFP19]. More families were studied later by many people; see [HP18, DP16, DMH21, LFP24, LFP19, LFP23]. In [LFP24], the authors gave the definition of semitorc families with degenerate times and semitoric transition families. A semitoric family with degenerate times $t_{1},\dots,t_{n}\in[0,1]$, is a family of integrable systems $(M,\omega,F_{t})$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that $(M,\omega,F_{t}:=(J,H_{t}))$, where $H_{t}$ is of the form $H_{t}=H(t,\cdot)$ for some smooth $H:[0,1]\times M\rightarrow{\mathbb{R}}$, is semitoric if and only if $t\in[0,1]\setminus\{t_{1},\dots,t_{n}\}$. In this paper, we study semitoric transition families on pentagon spaces.
A semitoric transition family with transition point $p\in M$ and transition times $t^{-},t^{+}$, with $0<t^{-}<t^{+}<1$, is a family of integrable systems $(M,\omega,F_{t}=(J,H_{t}))$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that:
@@
  • for $t\in[0,t^{-})\cup(t^{+},1]$, the singular point $p$ is of elliptic-elliptic type;
  • for $t\in(t^{-},t^{+})$, the singular point $p$ is of focus-focus type;
  • for $t=t^{-}$ and $t=t^{+}$, there are no degenerate singular point in $M\setminus\{p\}$, and
  • if $p$ is a maximum (respectively minimum) of $(H_{0})|_{J^{-1}(J(p))}$ then $p$ is a minimum (respectively maximum) of $(H_{1})|_{J^{-1}(J(p))}$.
@@
The $\mathbf{n}$-polygon space $\mathcal{P}(r_{1},\ldots,r_{n})$ is the quotient space $\mathcal{P}(r_{1},\ldots,r_{n}):=\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})/SO(3)$. In other words, the polygon space consists of the configurations in ${\mathbb{R}}^{3}$ of polygons with side lengths $r_{1},\ldots,r_{n}$, up to rotation. We say $r_{1},\ldots,r_{n}$ are generic if there does not exist a map $\epsilon:\{1,\ldots,n\}\to\{\pm 1\}$ such that $\sum_{i=1}^{n}\epsilon_{i}r_{i}=0$. Generic side lengths guarantee that the polygon space, if not empty, is a smooth manifold of dimension 2n62n-6. These spaces come with symplectic structures. Consider $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$, the product of spheres of radii $r_{1},\ldots,r_{n}$, with symplectic form $\frac{1}{r_{1}^{2}}\omega\oplus\cdots\oplus\frac{1}{r_{n}^{2}}\omega$, where $\omega_{u}(v,w)=u\cdot(v\times w)$. Here, we think of $S^{2}_{r_{i}}$ as a submanifold of ${\mathbb{R}}^{3}$. The map $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ given by $\mu(\rho_{1},\ldots,\rho_{n})=\sum_{i=1}^{n}\rho_{i}$ is the moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$, where we identify $\mathfrak{so}_{3}^{*}$ with ${\mathbb{R}}^{3}$ equipped with the cross product. Then the symplectic reduction at level 0 gives us $\mathcal{P}(r_{1},\ldots,r_{n})=\mu^{-1}(0)/SO(3)$, so it comes with a symplectic structure. See [HK97, Kly94, KM95, KM96]. Let $I\subset\{1,\ldots,n\}$ be a nonempty proper subset. Define $\ell_{I}:\mathcal{P}(r_{1},\ldots,r_{n})\to{\mathbb{R}}$ by $\ell_{I}([\rho])=|\sum_{i\in I}\rho_{i}|$.
@@
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $\frac{1}{2}r_{5}\leq r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$, and
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Then the family $F_{t}:=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ is a semitoric transition family on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$, with transition point $P$ represented by the planar pentagon, whose five edges are given by
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j^{2}}{2j},\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j^{2}}{2j},-\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{3}=(-r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(r_{5},0,0),$
and transition times
$t^{\pm}=\frac{\,r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j}}{\,(r_{3}+r_{5})^{2}+4r_{4}j\,}.$
Here $j=r_{3}+r_{4}-r_{5}$.
Theorem 1.1 gives new explicit examples of semitoric systems. The index sets in the statement are chosen for convenience. In fact, one may take any three distinct index sets $I,J,K$ of size $2$ such that $I$ is disjoint from both $J,K$ and consider the families $(\ell_{I},t\ell_{J}^{2}+(1-t)\ell_{K}^{2})$ on pentagon spaces whose side lengths satisfy similar constraints to those in Theorem 1.1. Moreover, different choices of side lengths $r_{1},\ldots,r_{5}$ that satisfy the assumptions of Theorem 1.1 give different symplectic forms on pentagon spaces of the same diffeotype, so these families will be useful in understanding how the change of symplectic form affects the transition time and other invariants of these semitoric systems. Additionally, our construction can be easily generalized to polygon spaces of higher dimensions to study integrable systems where some components of the functions give a moment map of a Hamiltonian torus action.
@@
In Section 2, we review some facts about polygon spaces and study an $S^{1}$-action on the pentagon space of interest. In Section 3, we review the basic concepts of semitoric systems and study the local theory of the transition point. In Section 4, we prove that the singular points other than the transition point $P$ are non-degenerate. In Section 5, we study the transition point $P$ and transition times. In Section 6 prove Theorem 1.1.

### v2

@@ Abstract.
Semitoric systems are a special type of 4-dimensional integrable system where one of the functions is the moment map of a Hamiltonian $S^{1}$-action. While their classification is well understood thanks to the work of Pelayo and Vũ Ngọc, relatively few explicit examples are known. Recently, Le Floch and Palmer introduced semitoric transition families in which a singular point transitions between elliptic-elliptic type and focus-focus type as the parameter varies. In this paper, we construct new semitoric transition families on pentagon spaces by interpolating two semitoric systems of toric type. More precisely, we exhibit semitoric transition families, each with at least one transition point, on pentagon spaces of different diffeotypes, all defined by the same explicit interpolation formula.
@@ 1. Introduction
An integrable system on a $2n$-dimensional symplectic manifold $(M,\omega)$ is a map $F=(f_{1},\ldots,f_{n}):M\to{\mathbb{R}}^{n}$ such that $\{f_{i},f_{j}\}=0$ for all $i,j$ and $df_{1},\ldots,df_{n}$ are linearly independent almost everywhere. A semitoric system is a special kind of integrable system when $\dim M=4$. More specifically, an integrable system $(M,\omega,F=(J,H))$ is a semitoric system if $J$ is a proper map that generates an effective Hamiltonian $S^{1}$-action on $M$ and the momentum map $F$ only has non-degenerate singularities without hyperbolic type (see Definition 3.3). Semitoric systems were introduced by  [VŨ 07] following the work of [SYM03]. Pelayo and Vũ Ngọc [PV09, PV11] first classified simple semitoric systems. Following their work, Pelayo-Palmer-Tang [PPT24] generalized the classification to any semitoric system. Semitoric systems have been intensively studied in recent years. See [ADH20, ADH19, AHP25, LP19] for studies of invariants of semitoric systems and [AH19, SV17, DKL+23, PAL25] for recent survey articles and additional references.
The coupled angular momenta system was the first example of a family of semitoric systems known to us. It was first introduced in [SZ99] and studied in detail in [LP19]. More families were studied later by many people; see [HP18, DP16, DH21, LP24, LP19, LP23]. In [LP24], the authors gave the definition of semitorc families with degenerate times and semitoric transition families. A semitoric family with degenerate times $t_{1},\dots,t_{n}\in[0,1]$, is a family of integrable systems $(M,\omega,F_{t})$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that $(M,\omega,F_{t}:=(J,H_{t}))$, where $H_{t}$ is of the form $H_{t}=H(t,\cdot)$ for some smooth $H:[0,1]\times M\rightarrow{\mathbb{R}}$, is semitoric if and only if $t\in[0,1]\setminus\{t_{1},\dots,t_{n}\}$. In this paper, we study semitoric transition families on pentagon spaces.
A semitoric transition family with transition point $p\in M$ and transition times $t^{-},t^{+}$, with $0<t^{-}<t^{+}<1$, is a family of integrable systems $(M,\omega,F_{t}=(J,H_{t}))$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that
@@
  • $p$ is a singular point of elliptic-elliptic type of $(M,\omega,F_{t})$ for $t\in[0,t^{-})\cup(t^{+},1]$;
  • $p$ is a singular point of focus-focus type of $(M,\omega,F_{t})$ for $t\in(t^{-},t^{+})$;
  • there are no degenerate singular points in $M\setminus\{p\}$ for $t=t^{\pm}$; and
  • if $p$ is a maximum (respectively, minimum) of $(H_{0})|_{J^{-1}(J(p))}$ then $p$ is a minimum (respectively, maximum) of $(H_{1})|_{J^{-1}(J(p))}$.
@@
The $\mathbf{n}$-polygon space $\mathcal{P}(r_{1},\ldots,r_{n})$ is the quotient space $\mathcal{P}(r_{1},\ldots,r_{n}):=\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})/SO(3)$. In other words, the polygon space consists of the configurations in ${\mathbb{R}}^{3}$ of polygons with side lengths $r_{1},\ldots,r_{n}$, up to rotation. We say $r_{1},\ldots,r_{n}$ are generic if there does not exist a map $\epsilon:\{1,\ldots,n\}\to\{\pm 1\}$ such that $\sum_{i=1}^{n}\epsilon_{i}r_{i}=0$. Generic side lengths guarantee that the polygon space, if not empty, is a smooth manifold of dimension 2n62n-6. These spaces come with symplectic structures. Consider $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$, the product of spheres of radii $r_{1},\ldots,r_{n}$, with symplectic form $\frac{1}{r_{1}^{2}}\omega\oplus\cdots\oplus\frac{1}{r_{n}^{2}}\omega$, where $\omega_{u}(v,w)=u\cdot(v\times w)$. Here, we think of $S^{2}_{r_{i}}$ as a submanifold of ${\mathbb{R}}^{3}$. The map $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ given by $\mu(\rho_{1},\ldots,\rho_{n})=\sum_{i=1}^{n}\rho_{i}$ is a moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$, where we identify $\mathfrak{so}_{3}^{*}$ with ${\mathbb{R}}^{3}$ equipped with the cross product. Then the symplectic reduction at the level 0 gives us $\mathcal{P}(r_{1},\ldots,r_{n})=\mu^{-1}(0)/SO(3)$, so it comes with a symplectic structure. See [HK97, KLY94, KM95, KM96]. Let $I\subset\{1,\ldots,n\}$ be a nonempty proper subset. Define $\ell_{I}:\mathcal{P}(r_{1},\ldots,r_{n})\to{\mathbb{R}}$ by $\ell_{I}([\rho])=|\sum_{i\in I}\rho_{i}|$.
Before stating our result, we first define two important pentagons $P,Q$ that will be transition points of $F_{t}$ and numbers $t_{P}^{\pm},t_{Q}^{\pm}$ that will be transition times. For positive numbers $r_{1},\ldots,r_{5}$, we define $j_{P}=r_{3}+r_{4}-r_{5}$, $j_{Q}=-r_{3}+r_{4}+r_{5}$. The pentagon $P\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is represented by the planar pentagon whose five edges are given by
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j_{P}^{2}}{2j_{P}},\frac{\sqrt{\big((r_{1}+j_{P})^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j_{P})^{2}\big)}}{2j_{P}},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j_{P}^{2}}{2j_{P}},-\frac{\sqrt{\big((r_{1}+j_{P})^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j_{P})^{2}\big)}}{2j_{P}},0\right),$
$\rho_{3}=(-r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(r_{5},0,0).$
The pentagon $Q\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is represented by the planar pentagon whose five edges are given by
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j_{Q}^{2}}{2j_{Q}},\frac{\sqrt{\big((r_{1}+j_{Q})^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j_{Q})^{2}\big)}}{2j_{Q}},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j_{Q}^{2}}{2j_{Q}},-\frac{\sqrt{\big((r_{1}+j_{Q})^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j_{Q})^{2}\big)}}{2j_{Q}},0\right),$
$\rho_{3}=(r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(-r_{5},0,0).$
The quantities $t_{P}^{\pm},t_{Q}^{\pm}$ are defined as follows.
$t_{P}^{\pm}=\frac{\,r_{4}j_{P}+r_{3}r_{5}+(r_{4}-r_{5})^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j_{P}}}{\,(r_{3}+r_{5})^{2}+4r_{4}j_{P}\,},\quad t_{Q}^{\pm}=\frac{\,r_{4}j_{Q}+r_{3}r_{5}+(r_{4}+r_{5})^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j_{Q}}}{\,(r_{3}+r_{5})^{2}+4r_{4}j_{Q}\,}.$
Notice that $P,Q$ are not always well-defined. More specifically, $P$ is well-defined in case (P),(P+Q), but not in case (Q) in Theorem 1.1 below and $Q$ is well-defined in case (Q),(P+Q), but not in case (P). Similarly, $t_{P}^{\pm}$ is a well-defined real number in case (P),(P+Q), but may not be a real number in case (Q) and vice versa.
@@
Let $r_{1},\ldots,r_{5}$ be positive numbers. Consider the family $F_{t}:=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$.
  1. (P)
    If $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$ and $r_{3}+r_{4}-r_{5}<r_{1}+r_{2}<r_{3}-r_{4}+r_{5}$, $F_{t}$ is a semitoric transition family with transition point $P$ and transition times $t_{P}^{\pm}$.
  2. (Q)
    If $|r_{3}+r_{4}-r_{5}|<|r_{1}-r_{2}|<-r_{3}+r_{4}+r_{5}$ and $r_{3}+r_{4}+r_{5}<r_{1}+r_{2}<r_{3}-r_{4}+r_{5}$, $F_{t}$ is a semitoric transition family with transition point $Q$ and transition times $t_{Q}^{\pm}$.
  3. (P+Q)
    If $r_{4}+|r_{3}-r_{5}|<|r_{1}-r_{2}|<r_{4}-|r_{3}-r_{5}|$, $r_{4}+|r_{3}-r_{5}|<r_{1}+r_{2}<r_{3}-r_{4}+r_{5}$ and $r_{3}\neq r_{5}$, $F_{t}$ is a semitoric transition family with two distinct transition points $P$ and $Q$ with transition times $t_{P}^{\pm}$ and $t_{Q}^{\pm}$ respectively.
Remark 1.2.
In case $(P+Q)$, we assume that $r_{3}\neq r_{5}$, which implies that $0<t_{P}^{-}<t_{P}^{+}<\frac{1}{2}<t_{Q}^{-}<t_{Q}^{+}<1$. When $r_{3}=r_{5}$, both $P$ and $Q$ are degenerate at the time $t_{P}^{+}=t_{Q}^{-}=\frac{1}{2}$, so the fourth bullet point in the definition of a semitoric transition family is not satisfied. In this case, $P$ and $Q$ are in the same fiber of $(J,H_{\frac{1}{2}})$ which is a Lagrangian two-sphere.
Proof of Theorem 1.1.
The proof of all three cases are almost identical, so we will only prove case (P). We write $F_{t}=(J,H_{t})$ and $M=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. By assumption, the moment images of $(J,\sqrt{H_{0}})$ and $(J,\sqrt{H_{1}})$ are both Delzant polygons (see Figure 3). Hence, for $t=0,1$, the system $F_{t}$ is of toric type. Moreover, $H_{1}(P)=(r_{3}+r_{4})^{2}$ is the global maximum of $H_{1}$ on $M$ and $H_{0}(P)=(r_{4}-r_{5})^{2}$ is the global minimum of $H_{0}$ on $M$. Hence, it follows from Proposition 5.1 that $P$ is a transition point with transition times $t_{P}^{\pm}$.
We now show that for all $t\in[0,1]$, all singular points of $F_{t}$ other than $P$ are non-degenerate of either elliptic-elliptic or elliptic-regular type. Fix a singular point $m\in M\setminus\{P\}$ and define $c:=J(m)$. If $c=|r_{1}-r_{2}|$ or $c=r_{1}+r_{2}$, then $J^{-1}(c)$ is a fixed surface by Lemma 2.5, so Proposition 2.2 implies that $J^{-1}(c)/S^{1}=J^{-1}(c)\cong\mathcal{P}(c,r_{3},r_{4},r_{5})$. Proposition 3.9 and Proposition 4.2 together imply that if $m$ is of rank one, then it is of elliptic-regular type, and if $m$ is of rank zero, then it is of elliptic-elliptic type.
Now assume that $c\neq|r_{1}-r_{2}|,r_{1}+r_{2}$. Since $m\neq P$, Lemma 2.5 implies that the $S^{1}$-action generated by $J$ acts freely on $m$ and thus $m$ is of rank one. Moreover, if we further assume $c\neq j_{P}$, then $c$ is a regular value of $J$ and $S^{1}$ acts freely on $J^{-1}(c)$. Hence, Proposition 2.2 implies that $M^{\textrm{red},c}:=J^{-1}(c)/S^{1}\cong\mathcal{P}(c,r_{3},r_{4},r_{5})$. Let $H_{t}^{\textrm{red},c}$ be the function on $M^{\textrm{red},c}$ to which $H_{t}$ descends, and $m]\in M^{\textrm{red},c}$ be the point to which $m$ descends. The first part of Lemma 3.7 implies that $[m]$ is a singular point of $H_{t}^{\textrm{red},c}$, which is an elliptic (in the Morse sense) singular point by Proposition 4.2. Hence, by the second part of Lemma 3.7, $m$ is an elliptic-regular singularity of $F_{t}$.
Finally if $c=j_{P}$, then $M^{\textrm{red},c}$ is not a manifold. However, since $S^{1}$ acts freely on $m$, there is an $S^{1}$-invariant neighborhood around the orbit of $m$ on which $S^{1}$ acts freely. Hence, one can do symplectic reduction locally, and verify that $[m]$ is an elliptic singularity of $H_{t}^{\textrm{red},c}$ by computing its Hessian in the same way as in the proof of Proposition 4.2. Therefore, any rank one points that are not on the fixed surfaces are of elliptic-regular type.
Therefore, we conclude that for $t\in[0,1]\setminus\{t_{P}^{-},t_{P}^{+}\}$, $(M,\omega,F_{t})$ is a semitoric system, and when $t=t_{P}^{\pm}$, there are no degenerate singular points in $M\setminus\{P\}$.
$\times$
(a) Case (P), $t\in(t_{P}^{-},t_{P}^{+})$
$\times$
(b) Case (Q), $t\in(t_{Q}^{-},t_{Q}^{+})$
[…]
(c) Case (P+Q), where $r_{3}>r_{5}$
[…]
(d) Case (P+Q), where $r_{3}<r_{5}$
Figure 1. Each figure is a presentative of the semitoric polygon of the corresponding case when the focus-focus point is present.
$\times$
$\times$
$\times$
$\times$
Figure 2. Other semitoric polygons that can be obtained from the same family $F_{t}$ by changing the inequalities between the parameters $r_{1},\dots,r_{5}$.
Remark 1.3.
Theorem 1.1 gives new explicit examples of semitoric systems. In this remark, we explain possible generalizations of our constructions and how it is related to other works.
  1. (1)
    The index sets in the statement are chosen for convenience. In fact, one may take any three distinct index sets $I,J,K$ of size $2$ such that $I$ is disjoint from both $J,K$ and consider the families $(\ell_{I},t\ell_{J}^{2}+(1-t)\ell_{K}^{2})$ on pentagon spaces whose side lengths satisfy constraints similar to those in Theorem 1.1.
  2. (2)
    Different choices of side lengths $r_{1},\ldots,r_{5}$ that satisfy the same set of inequalities give different symplectic forms on pentagon spaces of the same diffeotype, so these families will be useful in understanding how the change of symplectic form affects the transition time and other invariants of these semitoric systems.
  3. (3)
    Our construction can be generalized to $n$-polygon spaces $\mathcal{P}(r_{1},\ldots,r_{n})$ with $n>5$ to study higher-dimensional integrable systems where some components of the functions give a moment map of a Hamiltonian torus action.
  4. (4)
    The semitoric polygon invariant is one of the invariants of a semitoric system, analogous to the Delzant polytopes in the classification of symplectic toric manifolds. For the definition, we refer the reader to [VŨ 07] and [PV09, Section 4]. By [LP24, Lemma 3.14], a representative of the semitoric polygon (when the focus-focus point is present) of these families can be obtained by adding a marked point and a cut line to the semitoric polygon of the system $(J,H_{1})$. We display these images in Figure 1.
  5. (5)
    We can obtain additional semitoric polygons shown in Figure 2 from the same family $F_{t}$ by changing the inequalities between the parameters $r_{1},\dots,r_{5}$. But since the proof for the additional cases is essentially the same, we only list the three cases (P),(Q), and (P+Q) for simplicity. In fact, the other cases shown in Figure 2 also follow from [LP24, Theorem 1.7] as these polygons are obtained by performing a toric type blowup.
  6. (6)
    By [HK97], pentagon spaces are generically symplectic toric four-manifolds, and there are in total five diffeotypes ${\mathbb{C}}{\mathbb{P}}^{2},(S^{2}\times S^{2})\#k\overline{{\mathbb{C}}{\mathbb{P}}^{2}}$, $k=0,1,2,3$. Le Floch and Palmer proved that ${\mathbb{C}}{\mathbb{P}}^{2}$ does not admit a semitoric transition family (see [LP23, Proposition 5.2] and the discussion there). Le Floch and Pelayo studied the semitoric transition family given by the coupled angular momenta system on $S^{2}\times S^{2}$ in detail (see [LP19]). Our results provide new explicit families for diffeotypes $(S^{2}\times S^{2})\#k\overline{{\mathbb{C}}{\mathbb{P}}^{2}}$, $k=1,2,3$.
@@
In Section 2, we review some facts about polygon spaces and study an $S^{1}$-action on the pentagon space of interest. In Section 3, we review the basic concepts of semitoric systems and study the local theory of the transition point. In Section 4, we study some Morse functions on quadrilateral spaces. We apply the main result, Proposition 4.2, to functions induced by $H_{t}$ on the symplectic reduced spaces $J^{-1}(c)/S^{1}$ in the proof of Theorem 1.1. In Section 5, we study the transition point and transition times.